A rectangular auditorium seats
1575
people. The number of seats in each row exceeds the number of rows by
10
. Find the number of seats in each row.

Answer by ikleyn(52787)

.
sqrt(1575) =~ 40.

Answer. 35 and 45. The number of seats is 45.



If you want more accurate math solution, here it is.

Let the unknown x be the number of rows PLUS 5.

Then the number of seats is (x-5).

According to the condition,  

(x+5)*(x-5) = 1575,

x^2 - 25 = 1575 --->  x^2 = 1575 + 25 = 1600  --->  x = 40.

Then you get the answer.

In this way you get the solution of the problem mentally, without using the quadratic formula.


See the lesson
    - HOW TO solve the problem on quadratic equation mentally and avoid boring calculations
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Quadratic equations".